# Dynamic Input Mapping Inversion for Algebraic Loop-Free Control in Hydraulic Actuators

### v2

@@ Abstract
The application of nonlinear control schemes to electro-hydraulic actuators often requires several alterations in the design of the controllers during their implementation. This is to overcome the challenges that frequently arise from the inherent complexity of such control algorithms owning to model nonlinearities. Moreover, advanced control solutions for this type of systems often introduce input algebraic loops and chatter, which considerably degrade the tracking performance. This study presents a nonlinear control architecture for hydraulic actuators that comprises low-complexity modules, based on well-established designs that facilitate robust high performance in tracking without introducing the aforementioned limitations. Specifically, the proposed solution consists of two variants of a position controller for the hydraulic cylinder and a dynamic input-mapping inversion module to avoid algebraic loops in the control input. The stability of the closed-loop system is analysed using arguments from Lyapunov theory for cascaded non-autonomous nonlinear systems. The effectiveness of the proposed solution is evaluated on a high-fidelity simulator of a wind turbine pitch system. Appropriate quantitative metrics are finally defined to evaluate the closed-loop system performance in comparison to state-of-the-art nonlinear design.
Index Terms:
electro-hydraulic actuator, nonlinear control, parameter estimation, mapping inversion
@@ I Introduction
Electro-hydrualic actuators (EHA) are commonly used in industrial applications because of their high power density and resilience to harsh environment. Typical applications include aviation, manufacturing, marine/offshore engineering, robotics and energy systems [1, 2, 3].
@@
Although Proportional-Integral-Differential (PID) controllers constitute the industrial standard in these application domains, they often fall short in maintaining high performance in the presence of strong nonlinearities and uncertainties that characterize hydraulic actuator systems. This has motivated the exploration of nonlinear control alternatives that can improve performance and enhance closed-loop robustness over the entire envelope of operation. The complexity, however, of system dynamics often hinders effective application of nonlinear model-based solutions such as feedback linearization (FL) or sliding mode control (SMC) because several limiting assumptions and modelling simplifications need be imposed to apply these methods.
@@
A common heavy simplification concerns the valve dynamics, where two main issues can be identified. A first limitation is encountered when a system model cannot be written as an affine form $\dot{x}=f(x)+g(x)u$, as the vector field of a valve depends non-linearly on the input. A common approach to address this issue is to use the input value at the previous time instant, under the assumption that the input $u$ is slowly varying. This removes the nonlinearity and avoids the introduction of an algebraic loop, at the cost of increased chatter in the control signal [4]. The second design challenge arises with inclusion of actuator dynamics, which increase the relative degree of the system, thus making the design much more complicated, as will be discussed later. Neglecting the valve dynamics during the design stage can simplify implementation but may potentially result in a quite demanding control action going beyond obtainable valve response. This may introduce additional chatter and potential wind up effects.
@@
Furthermore, micro-movements in EHAs are responsible for local heating of hydraulic oil, which degrades oil quality. This leads to increased friction, and eventually to wear and potential damage of a valve [5]. It is therefore essential to establish control designs that combine the performance and robustness features of nonlinear solutions, but avoid excessive actuation demand.
@@
Several studies in the existing literature have reported solutions that adopt nonlinear strategies for controlling hydraulic systems. A cascaded feedback linearization controller with a sliding mode observer was designed in [6]. Sliding mode was explored in a standard fashion in [7], in combination with adaptive control in [8], with an integral sliding surface in [9] and later in [10] to eliminate the velocity state in the control action. Adaptive solutions have been applied by means of backstepping, in e.g. [11], [12], [13], and [14]. The authors in [15] used a coordinate transformation and tuning functions to obtain a backstepping controller that included the valve dynamics. The high relative degree led to a complex solution. Except for the latter, all the other solution adopted the strategy presented in [4] to cancel the algebraic loop, and do not account for the valve dynamics in the design of the controller.
@@
The aim of this paper is to leverage the results on online estimation of nonlinearly parametrised perturbations [16, 17] as a dynamical mapping inversion (MI) algorithm to derive the control input from a virtual input designed via classical nonlinear methods. The paper shows how a modular and straightforward design is achieved, and how the MI can reconstruct the control signal without introducing algebraic loops. Moreover, it is shown how the suggested approach facilitates a systematic commissioning feature, namely, how it can accommodate valve specification parameters such that only frequencies that can actually be delivered by the valve are allowed in the control input. In this way, the performance can be improved and chatter in the input be significantly reduced.
This article contributes by
@@
    presenting a modular architecture for nonlinear control of hydraulic actuators without introducing algebraic loops or limiting assumptions on the input signal,
@@
    proposing a scheme for reconstructing the control signal by leveraging a nonlinear parameter estimation algorithm for dynamically inverting the mapping between control input and virtual input,
@@
    introducing a systematic method to tune the dynamical mapping inversion component according to the actuator parameters,
  4. 4.
    providing rigorous stability and boundedness analyses of the closed-loop dynamics under different control laws,
  5. 5.
    demonstrating the effectiveness of the proposed solution against conventional nonlinear control methods for industrial EHA systems in a real case scenario, using a high-fidelity simulator.
The remainder of the paper is organized as follows. Sec. II presents the hydraulic actuator system and defines the problem at hand. Sec. III designs controllers following commonly accepted standards, and Sec. IV proposes a new algorithm and presents features and benefits in contrast to the standard existing approaches. Stability of the closed-loop system is the subject of Sec. V. Sec. VI presents a case study of a wind turbine pitch control, and results from high fidelity simulation are shown in Sec. VII. Finally, conclusions and future directions are staged in Sec. VIII.
@@ II Problem definition
An electro-hydraulic actuator is composed of an hydraulic cylinder driven by a directional control valve, as shown in Fig. 1. The pressures $p_{p}$ and $p_{r}$ in the piston-side and rod-side chambers of the cylinder are varied according to the valve spool position $x_{v}$. This allows for control of both position and velocity ($x_{c}$ and $v_{c}$) of the the cylinder’s piston. The model for a generic electro-hydraulic actuator is given by [9]:
@@
Figure 1: Generic electro-hydraulic actuator configuration: above, the linear cylinder with piston and rod; below, the flow control valve, oil pump and drain.
where $M$ is the piston mass, $A_{p}$ and $A_{r}$ the cylinder cross sections in the two chambers, $F_{fr}$ and $F_{ext}$ the friction force and the external load respectively, and $B_{e,i}(p_{i})$ the effective bulk modulus of the oil in the $i$-th component. The control input $u_{x}$ is a position reference for the control valve’s spool. The terms
@@
with $H(\cdot)$ being the Heaviside step function.
The control objective in EHA systems is the tracking of a reference signal for the cylinder piston position $x_{c}$. By inspecting the cylinder position dynamics in Eq. 1 it is easily seen that the system corresponding to the position error defined as
$e\triangleq x_{c}-r,$ (5)
has relative degree 5, i.e. $e$ needs to be differentiated five times in order to arrive to a controller canonical form
@@
For the class of EHA systems this introduces a high degree of complexity due to the strong nonlinear couplings between the various pressure terms. A common approach adopted in the literature to alleviate this complexity is the introduction of a simplifying assumption, namely the neglection of the valve dynamics. This implies that the position $x_{v}$ inside the terms $Q_{p}(x_{v})$ and $Q_{r}(x_{v})$ can instantaneously assume any desired value and is, therefore, considered the effective control input for the system, i.e. $u\triangleq x_{v}$, with $u\in[-1,1]$.
An additional frequent simplification made in many reported studies is considering one single constant value for the bulk modulus of the oil for both chambers, i.e., $B_{e,p}=B_{e,r}=B>0$. The resulting model, which is most commonly used in the industry for control design, is [18]:
@@ III Baseline nonlinear control design
Considering the model presented in Eq. 7 and the position error $e$ defined in Eq. 5, the dynamics constitutes a relative degree 3 system and reads:
@@
$\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u)\triangleq\,\varpi_{FL}.$
@@
$\ddot{e}$ $=-\frac{F_{fric}+F_{ext}}{M}+\frac{A_{p}p_{p}-A_{r}p_{r}}{M}-\ddot{r},$
where the virtual input $\varpi_{FL}=\varpi_{FL}(t,\boldsymbol{x},u)$ is a function of time, of states $\boldsymbol{x}\triangleq\begin{bmatrix}v_{c}&x_{c}&F_{h}\end{bmatrix}^{T}$, and of the control input $u$.
Two basic control solutions are first presented to demonstrate both the effectiveness and the reasons behind the development of a new algorithm. Feedback linearization is presented as a baseline nonlinear control design, which is popular for adaptive nonlinear control solutions. A supertwisting sliding mode control design is shown as a second popular solution.
III-A Feedback linearization (FL)
@@
A FL controller is obtained from writing the error derivative dynamics as follows,
@@
$\dot{\underline{\mathbf{e}}}=\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix}}_{\boldsymbol{A_{FL}}}\underbrace{\begin{bmatrix}e\\ \dot{e}\\ \ddot{e}\end{bmatrix}}_{\underline{\mathbf{e}}}+\underbrace{\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}}_{\boldsymbol{B_{FL}}}\varpi_{FL}.$ (10)
Selecting
$\varpi_{FL}=-K_{FL}\underline{\mathbf{e}}$ (11)
@@
with $K_{FL}$ obtained with a standard linear design (eignvalue assignment, LQR, etc.), renders the closed-loop system globally exponentially stable (GES) with the rate of decay being defined by the gain $K_{FL}$. The related control input $u$ is isolated from Eq. 11 taking also Eq. 8, and Eq. 9 into account.
III-B Supertwisting SMC
Sliding mode controllers aim at enforcing the trajectories $\boldsymbol{e}(t)\in\mathbb{R}^{3}$ of the error system onto a pre-designed manifold $\mathcal{S}=\{\boldsymbol{e}\in\mathbb{R}^{3}\,\big{|}\,s(\boldsymbol{e})=0\}$. The variable $s$ is referred to as the sliding variable and is designed to guarantee exponential stability of the error system, so long $s$ is regulated at the origin.
In order to leverage the reduced chatter properties of the well-studied supertwisting sliding mode control designs [19], the dynamics of s should be brought into the form of Eq. 6 with a single differentiation. Therefore, the sliding variable $s$ is defined as:
$s=\ddot{e}+\lambda_{2}\dot{e}+\lambda_{1}e,$ (12)
where $\lambda_{1},\lambda_{2}$ are selected such that the characteristic polynomial in Eq. 12 is Hurwitz. The dynamics of the sliding variable $s$ is:
$\dot{s}=\underbrace{e^{(3)}+\lambda_{2}\ddot{e}+\lambda_{2}\dot{e}}_{\displaystyle\varpi_{ST}(t,\boldsymbol{x},u)},$ (13)
where the right-hand side is of Eq. 13 is again a function $\varpi_{ST}$ of time, the states and the control input $u$. Eq. 12 implies global exponential stability for the origin of the piston position error when the system is in sliding mode, i.e. when $s=\dot{s}=0$. This can be achieved by selecting
$\varpi_{ST}=-k_{1}\lfloor s\rceil^{1/2}-k_{2}\int_{0}^{t}sgn(s(\tau))d\tau,$ (14)
where $k_{1},k_{2}>0$. Similar to the controller case in the previous subsection, the control input is derived by isolating the term $u$ in equation 14.
III-C Control input
The final step for designing the controllers is deriving the control input $u$ for which the closed-loop system dynamics is shaped, via the previously presented design methods. From Eq. 8 it is noticed that the error dynamics depends on the terms $F_{ext}$ and $F_{fr}$, and their derivatives. Such terms are in general unknown and need be estimated to compute the control input $u$.
III-C1 Disturbance observer
Although dedicated tests can be run to identify the friction coefficients, it is convenient in this investigation to assume the friction term to be unknown. This allows for lumping it together with the external load in a single term $F_{d}=F_{ext}+F_{fr}$, which can then be estimated with a conventional linear observer. Such an observer can be designed based on the measured pressures as follows,
$\dot{\hat{\boldsymbol{z}}}=\underbrace{\begin{bmatrix}0&-\frac{1}{M}\\ 0&0\end{bmatrix}}_{\boldsymbol{A}}\hat{\boldsymbol{z}}+\begin{bmatrix}\displaystyle\frac{A_{p}p_{p}-A_{r}p_{r}}{M}\\ 0\end{bmatrix}+\underbrace{\begin{bmatrix}L_{1}\\ L_{2}\end{bmatrix}}_{\boldsymbol{L}}\underbrace{\begin{bmatrix}1&0\end{bmatrix}}_{\boldsymbol{C}}\tilde{\boldsymbol{z}},$ (15)
where $\hat{\boldsymbol{z}}=\begin{bmatrix}\hat{v_{c}}&\hat{F}_{d}\end{bmatrix}^{T}$, $\tilde{\boldsymbol{z}}\triangleq\boldsymbol{z}-\hat{\boldsymbol{z}}$ and $\boldsymbol{L}$ chosen such that $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz, ensuring that the disturbance term $F_{d}$ can be reconstructed with a bounded error [20]. Furthermore, since $\boldsymbol{z}\in\mathcal{L}_{\infty}$ and $\boldsymbol{A}-\boldsymbol{LC}$ is Hurwitz by design, it holds that
$\|\boldsymbol{\dot{\tilde{z}}}\|=\|(\boldsymbol{A}-\boldsymbol{LC})\,\boldsymbol{\tilde{z}}\|\leq-\lambda_{min}(\boldsymbol{A}-\boldsymbol{LC})\,\|\boldsymbol{\tilde{z}}\|_{\infty}\triangleq L_{z},$ (16)
where $L_{z}$ is a positive real number. Thus, $\dot{F}_{d}$ is also reconstructed with a bounded error. Global asymptotic stability of the entire closed-loop system will be proved subsequently.
III-C2 Control input derivation

### v4

# Dynamic Input Mapping Inversion to Eliminate Algebraic Loops in Hydraulic Actuator Control

@@ Abstract
The application of nonlinear control schemes to electro-hydraulic actuators often requires several alterations in the design of the controllers during their implementation. This is to overcome challenges that frequently arise in such control algorithms owing to model nonlinearities. Moreover, advanced control solutions for this type of systems often introduce input algebraic loops that pose significant design and tuning difficulties. Conventional methods to avoid such loops introduce chatter, which considerably degrade tracking performance and has oil degradation and wear as side effects. This study presents a nonlinear control architecture for hydraulic actuators that comprises low-complexity modules that facilitate robust high performance in tracking and avoids the drawbacks of chatter. The salient feature is a dynamic input-mapping inversion module that avoids algebraic loops in the control input and is followed by dedicated position control. The stability of the closed-loop system is analyzed using arguments from Lyapunov theory for cascaded non-autonomous nonlinear systems. The effectiveness of the proposed solution is evaluated on a high-fidelity simulator of a wind turbine pitch system, and validated on a full-scale laboratory setup that includes a hydraulic pitch system and blade bearing. Appropriate quantitative metrics are used to evaluate the closed-loop system performance in comparison to a state-of-the-art nonlinear design.
@@ I Introduction
Electro-hydrualic actuators (EHA) are commonly used in industrial applications because of their high power density and resilience to harsh environment. Typical applications include aviation, manufacturing, marine/offshore engineering, robotics and energy systems [fossenMarineControlSystems2002, walgernReliabilityElectricalHydraulic2023, dallabonaFaultDiagnosisPrognosis2025].
@@
Although Proportional-Integral-Differential (PID) controllers constitute the industrial standard in these application domains, they often fall short in maintaining high performance in the presence of strong nonlinearities and uncertainties that characterize hydraulic actuator systems. This has motivated the exploration of nonlinear control alternatives that can improve performance and enhance closed-loop robustness over the entire envelope of operation. The complexity, however, of system dynamics often hinders effective application of nonlinear model-based solutions because several limiting assumptions and modeling simplifications need be imposed to apply these methods.
@@
A common heavy simplification concerns the valve dynamics, where two main issues can be identified. A first limitation is encountered when a system model cannot be written in the affine form $\dot{x}=f(x)+g(x)u$, as the vector field depends non-linearly on the input. A common approach to address this issue is to use the input value at the previous time instant, under the assumption that the input $u$ is slowly varying. This removes the nonlinearity and avoids the introduction of an algebraic loop, at the cost of increased chatter in the control signal [ayalewCascadeTuningNonlinear2006]. The second design challenge arises with inclusion of actuator dynamics, which increases the relative degree of the system, thus making the design much more complicated, as will be discussed later. Neglecting the valve dynamics during the design stage can simplify implementation but may potentially result in a quite demanding control action going beyond obtainable valve response. This may introduce additional chatter and potential wind up effects.
@@
Furthermore, micro-movements in EHAs are responsible for local heating of hydraulic oil, which degrades oil quality. This leads to increased friction, and eventually to wear and potential damage of the valve [dallabonaFrictionEstimationCondition2024a]. It is therefore essential to establish control designs that combine the performance and robustness features of nonlinear solutions, but avoid excessive actuation demand.
@@
Several studies in existing literature have reported solutions that adopt nonlinear strategies for controlling hydraulic systems. A cascaded feedback linearization controller with a sliding mode observer was designed in [kochObserverbasedSlidingMode2016]. Sliding mode was explored in a standard fashion in [thomasPracticalProfileTracking2022], [guanAdaptiveSlidingMode2008a], for high order sliding modes in [estradaHydraulicActuatorControl2025] and in [estradaSupertwistingBasedSliding2024] to eliminate the velocity state in the control action. The authors in [yamamotoForceControllerValvemanipulated2025] and [yamamotoPositionControllerHydraulic2025] combined sliding mode with a quasi-static EHA model using a set-valued formalism to capture model nonlinearities. A fixed-time Event-Triggered Control strategy has been tested in [wangDisturbanceObserverBasedFixedTime2025]. Adaptive solutions have been applied by means of backstepping, in e.g. [chengguanNonlinearAdaptiveRobust2008], [chouxCascadeControllerIncluding2012], [bakhshandeProportionalIntegralObserverBasedBacksteppingApproach2018], [manganasBacksteppingBasedController2023], and neural networks in [guoNeuralAdaptiveBackstepping2019], [yangNeuralAdaptiveDynamic2023], [phanAdaptiveNeuralObserverbased2025]. The authors in [chouxAdaptiveBacksteppingControl2010a] used a coordinate transformation and tuning functions to obtain a backstepping controller that included the valve dynamics. The high relative degree led to a complex solution that would also be affected by valve degradation over time.
Except for said article, all other published solutions adopted the strategy presented in [ayalewCascadeTuningNonlinear2006] to cancel the algebraic loop, and did not account for the input valve dynamics in the design of the controller.
@@
This paper uses results from online estimation of nonlinearly parametrized perturbations [gripParameterEstimationCompensation2010b, papageorgiouRobustBacklashEstimation2019] to obtain a dynamical mapping inversion (MI) algorithm and derive the control input from a virtual input designed via classical nonlinear methods, without relying on the simplifying assumptions commonly adopted in the literature to implement nonlinear designs. The paper shows how a modular and straightforward design is achieved, and how the MI can reconstruct the control signal without introducing algebraic loops. Moreover, it is shown how the suggested approach facilitates a systematic commissioning feature, namely, how it can accommodate valve specification parameters such that only frequencies that can actually be delivered by the valve are allowed in the control input. In this way, the performance can be improved and chatter in the input can be significantly reduced.
The contributions of this article are:
@@
    A dynamical input mapping inversion algorithm in a modular architecture for control of hydraulic actuators that avoids algebraic loops and significantly reduces EHA micro-movements and associated wear.
@@
    A systematic method to tune the dynamical mapping inversion component according to control valve dynamics.
@@
    Rigorous analysis of stability/boundedness and validation on a wind-turbine EHA system, using both a high-fidelity simulator and a full-scale physical test rig.
The remainder of the paper is organized as follows. Sec. II presents the hydraulic actuator system and defines the problem at hand. Sec. III designs a well-established baseline controller following commonly accepted standards, and Sec. IV proposes a new algorithm and presents features and benefits in contrast to the standard existing approaches. Stability of the closed-loop system is the subject of Sec. V. Sec. VI presents a case study of a wind turbine pitch control, and results from validation on a full-scale electro-hydraulic test rig are shown in Sec. VII. Finally, conclusions and future directions are offered in Sec. VIII.
@@ II Problem definition
An electro-hydraulic actuator is composed of an hydraulic cylinder driven by a directional control valve (DCV), as shown in Fig. 1. The pressures $p_{p}$ and $p_{r}$ in the piston-side and rod-side chambers of the cylinder are varied according to the valve’s spool position $x_{v}$. This allows for control of both position and velocity ($x_{c}$ and $v_{c}$) of the the cylinder’s piston. The model for a generic electro-hydraulic actuator is given by the dynamic model 1, see e.g. [komstaIntegralSlidingMode2013, dallabonaFaultDiagnosisPrognosis2025, pedersenInvestigationLoadReduction2016]:
@@
Figure 1: Generic electro-hydraulic actuator system: above, the linear cylinder with piston and rod; below, the flow control valve, oil pump and drain.
where $M$ is the piston mass, $A_{p}$ and $A_{r}$ the cylinder cross sections in the two chambers, $F_{fr}$ and $F_{ext}$ the friction force and the external load respectively, and $B_{e,i}(p_{i})$ the effective bulk modulus of the oil in the $i$-th component. The terms
@@
where $H(\cdot)$ stands for the Heaviside step function, and $p_{s}$ and $p_{t}$ for system and tank pressures, respectively.
The terms $\omega_{0}$ and $\xi$ denote the natural frequency and damping ratio of the input valve, which is modeled as a second-order system. This stems from DCVs typically being equipped with an internal closed loop. As a result, the control input $u_{x}$ is a normalized, thus dimensionless, position reference for the valve spool.
The control objective in EHA systems is the tracking of a reference signal for the cylinder piston position $x_{c}$. By inspecting the cylinder position dynamics in (1) it is easily seen that the system corresponding to the position tracking error $e$,
$e=x_{c}-r,$ (5)
has relative degree $\alpha=5$, i.e. $e$ needs to be differentiated five times in order to arrive to a controller canonical form
@@
For the class of EHA systems this introduces a high degree of complexity due to the strong nonlinear couplings between the various pressure terms. A common approach adopted in the literature to alleviate this complexity is the introduction of a simplifying assumption, namely neglecting the valve dynamics. This implies that the position $x_{v}$ inside the terms $Q_{p}(x_{v})$ and $Q_{r}(x_{v})$ can instantaneously assume any desired value and is, therefore, considered the effective control input for the system, i.e. $u\triangleq x_{v}$ and $Q_{p}(x_{v}),Q_{r}(x_{v})$ becoming $Q_{p}(u),Q_{r}(u)$.
An additional frequent simplification made in many reported studies is considering one single constant value for the bulk modulus of the oil for both chambers, i.e., $B_{e,p}=B_{e,r}=B>0$. The resulting model, which is most commonly used in the industry for control design, is [bakhshandeRobustControlApproach2017]:
@@ III Baseline nonlinear control design
The control architecture for the baseline solution is presented in Fig. 2, and represent the standard configuration for nonlinear control solutions for the system under analysis.
Considering the model presented in (7) and the tracking error $e$ defined in (5), the dynamics constitutes a relative degree 3 system and reads,
@@
$\frac{A_{p}B}{MV_{p}}Q_{p}(u)+\frac{A_{r}B}{MV_{r}}Q_{r}(u)\triangleq\,\varpi.$
where the term $\varpi=\varpi(t,\boldsymbol{x},u)$ is a function of time, of states $\boldsymbol{x}\triangleq\begin{bmatrix}v_{c}&x_{c}&F_{h}\end{bmatrix}^{T}$, and of the control input $u$.
@@
$\ddot{e}$ $=-\frac{F_{fr}+F_{ext}}{M}+\frac{A_{p}p_{p}-A_{r}p_{r}}{M}-\ddot{r},$
Figure 2: Block diagram of the baseline algorithm that does not consider DCV dynamics.
Feedback linearization (FL) is chosen as the baseline nonlinear control design to showcase the benefits of the new developed algorithm, given its widespread adoption in adaptive control schemes. Other solutions like sliding mode control, which often rely on an equivalent control law for satisfactory performance[dallabonaModelfreeSlidingMode2024], could similarly be used.
@@
A feedback linearization (FL) control is obtained from writing the error derivative dynamics as follows,
@@
$\dot{\underline{\mathbf{e}}}=\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix}}_{\boldsymbol{A_{FL}}}\underbrace{\begin{bmatrix}e\\ \dot{e}\\ \ddot{e}\end{bmatrix}}_{\underline{\mathbf{e}}}+\underbrace{\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}}_{\boldsymbol{B_{FL}}}\varpi.$ (10)
Selecting $\varpi$ as
$\varpi=-K_{FL}\underline{\mathbf{e}},$ (11)
@@
with $K_{FL}$ obtained by a standard linear design (eigenvalue assignment, LQR, or other), renders the closed-loop system globally exponentially stable (GES) with the rate of decay being defined by the gain $K_{FL}$. The related control input $u$ is isolated from (11) taking also (8), and (9) into account.
From the latter it is noticed that the error dynamics depends on the terms $F_{ext}$ and $F_{fr}$, and their derivatives. Such terms are in general unknown and need be estimated to later compute the control input $u$. Although dedicated offline tests could identify the friction coefficients as they are during commissioning, the friction will change over time and it is both convenient and necessary to assume friction to be unknown. This allows for lumping friction together with external load as a single term $F_{d}=F_{ext}+F_{fr}$, which can be estimated by a conventional observer, using pressure measurements. One possible solution to this estimation problem was presented in [dallabonaModelfreeSlidingMode2024].
III-A Standard control input derivation